### v1

@@ 7 A simpler proof of the theorem of Snoeyink and Hershberger
We first need some terminology. Let $R$ be a digon bounded by the sweep curve $\gamma$ and another curve $\mu$ with vertices $b_{1}$ and $b_{2}$. A proper chord $\lambda$ of $R$ is an arc of another curve $\delta$ which intersects the boundary of $R$ exactly once on $\mu$ and once on $\gamma$. We say that a sequence of proper chords $\lambda_{1},\cdots,\lambda_{k}$ are parallel if for $i=1,\ldots,k-1$, the region of $R$ between consecutive chords $\lambda_{i}$ and $\lambda_{i+1}$ is a four-sided cell in the arrangement that does not contain $P$. For each $i$, let $\delta_{i}$ be the curve containing $\lambda_{i}$ and let $\Delta_{\mu}$ denote the sequence $\Delta_{\mu}=(\delta_{1},\ldots,\delta_{k})$.
@@
It is possible to modify the curves in $\Delta_{\mu}=(\delta_{1},\cdots,\delta_{k})$, possibly discarding some of them, so that i) the modified arrangement is as simple, or simpler than $\Gamma$, ii) the new digon/triangle cells created in the modified arrangement in $\tilde{\gamma}$ have either $b_{1}$ or $b_{2}$ as a vertex, and iii) there are no sweepable digon/triangle cells in $\tilde{\gamma}$ on $\mu$ other than those having $b_{1}$ or $b_{2}$ as one of the corners.
@@
If there are no digon/triangle cells on $\beta$, then there is nothing to do. Suppose there is a digon cell on $\beta$ formed by curve $\delta_{i}$. Then, $\Gamma\setminus\{\delta_{i}\}$ is a simpler arrangement with fewer digon/triangle cells on $\beta$. Since $\delta_{i}$ forms a digon cell on $\beta$, it forms a side of a cell $C$ in $\tilde{\gamma}\setminus R$ on $\beta$. Removing $\delta_{i}$ therefore decreases the number of sides in $C$ by 1, and does not affect any other cell in $\tilde{\gamma}\setminus R$. Hence, only $C$ can become a triangle. If $\delta_{i}$ and $\delta_{i+1}$ form a triangle cell $T$ on $\beta$, let $p$ be the vertex defined by $\delta_{i}$ and $\delta_{i+1}$ that forms a vertex of $T$. Then, the triangle formed by $\delta_{i}$ and $\delta_{i+1}$ on $\gamma$ is a minimal triangle with base $\gamma$. Hence, by Corollary 5.4 we obtain a simpler 2-intersecting arrangement. Since $T$ is intersected by $\beta$, by Lemma 7, the only cells with reference points $p_{\delta_{i}}$ or $p_{\delta_{i+1}}$ can become triangle cells. Therefore, the number of triangle cells on $\beta$ increases by 2, and no new triangle cells are created. If $\delta_{i}$ is not the first or last curve in the sequence $\Delta$, then consider the two cells with reference points $p_{\delta_{i}}$ or $p_{\delta_{i+1}}$. If either cell is a cell on $\gamma$, then they have at least 4 sides, as they have one of $\delta_{i}$ or $\delta_{i+1}$, $\beta$, and one of $\delta_{i^{\prime}}$ as three sides, besides $\gamma$ as sides, where $i^{\prime}<i$ or $i^{\prime}>i+1$.
At each iteration, we obtain a simpler arrangement where either the number of curves in $\Delta$ decreases, or the number of intersections between the curves $\Delta$ decreases. Therefore, after a finite number of iterations, there are no triangle cells or digon cells on $\beta$ except possibly the ones with $b_{1}$ or $b_{2}$ as a vertex. Figures 15 show the modification of the curves in $\Delta_{\mu}$.
@@
Handling triangle cells $C_{\alpha}$ and/or $C_{\beta}$: Suppose that $C_{\alpha}$ is a triangle cell $T$ that does not contain $P$, defined by the curves $\alpha,\gamma$ and $\delta_{1}$. Let $R$ be the digon defined by $\alpha$ and $\gamma$ that contains $T$. Without loss of generality let $a_{1},c_{1},d_{1}$ denote the three vertices of $T$, where $c_{1}$ and $d_{1}$ are the intersection points of $\delta_{1}$ with $\gamma$ and $\alpha$, respectively. Let $\Delta_{\alpha}=(\delta_{1},\ldots,\delta_{k})$ be a maximal sequence of parallel curves in $R$ starting with $\delta_{1}$. Since Let $d_{k}$ be the intersection point of $\delta_{k}$ and $\gamma$ in $R$. Since $\Gamma$ is a non-sweepable arrangement, either the arc of $\gamma$ between $d_{k}$ to $b_{2}$ does not bound a triangle cell, or the triangle cell contains $P$, and is therefore non-sweepable. Hence, the arc of $\gamma$ between $d_{k}$ to $b_{2}$ is intersected by a curve in $\Gamma\setminus\{\beta,\delta_{k}\}$. Similarly, since $\Gamma$ is non-sweepable, the arc of $\gamma$ between $a_{1}$ and $b_{1}$ is intersected by a curve in $\Gamma\setminus\{\beta,\gamma\}$, or is a triangle cell containing $P$. We first apply Lemma 7.6 with $\mu=\alpha$, $R$, and $\Delta_{\alpha}$. Let $\Delta_{\alpha}=(\delta^{\prime}_{1},\ldots,\delta^{\prime}_{\ell})$ denote the sequence of parallel curves after the application of Lemma 7.6. By Lemma 7.6, the possible sweepable triangle cells in the new arrangement are those in $\tilde{\gamma}$ that contain $a_{1}$ or $a_{2}$ as a vertex, i.e., the cells containing reference points $x_{1},y_{1},x_{2}$ or $y_{2}$ (See Figure 15(a)). The cell containing reference point $y_{1}$ cannot be a sweepable cell it either contains $P$, or there is at least one more curve other than $\gamma$ defining this cell that does not intersect $\alpha$ in $\tilde{\gamma}$. Similarly, the cell containing the reference point $x_{2}$ cannot be a sweepable cell, as this would imply that the cell in $\Gamma$ containing $x_{2}$ was already sweepable. Therefore, only the cells containing reference points $x_{1}$ or $y_{2}$ can be sweepable triangle cells. Observe that $\alpha,\gamma$ and $\delta^{\prime}_{\ell}$ form a minimal triangle with base $\delta^{\prime}_{\ell}$. By Corollary 5.4, bypassing this minimal triangle yields a simpler non-sweepable arrangement of 2-intersecting curves such that the two cells with reference points $x_{1}$ and $y_{1}$ merge to create a non-sweepable cell, and the cell containing reference point $y_{2}$ becomes a 4-sided cell. See Figure 15(b). Thus, we obtain a simpler non-sweepable arrangement. If $C_{\beta}$ is a triangle cell, we treat it similarly. Thus, in all cases, we obtain a simpler non-sweepable arrangement, contradicting the assumption that $\Gamma$ was a simplest non-sweepable arrangement. Hence, we can apply one of the sweep operations, and by Corollary 3, each sweep operation preserves the 2-intersecting property of the curves.

### v2

@@ 7 A simpler proof of the theorem of Snoeyink and Hershberger
We start with some necessary terminology. Let $R$ be a digon bounded by the sweep curve $\gamma$ and another curve $\mu$. A proper chord $\lambda$ of $R$ is an arc of another curve $\delta$ which intersects the boundary of $R$ exactly once on $\mu$ and once on $\gamma$. We say that a sequence of proper chords $\lambda_{1},\cdots,\lambda_{k}$ are parallel if for $i=1,\ldots,k-1$, the region of $R$ between consecutive chords $\lambda_{i}$ and $\lambda_{i+1}$ is a four-sided cell bounded by $\gamma,\lambda_{i},\mu$ and $\lambda_{i+1}$ in the arrangement that does not contain $P$. For each $i$, let $\delta_{i}$ be the curve containing $\lambda_{i}$ and let $\Delta_{\mu}$ denote the sequence $\Delta_{\mu}=(\delta_{1},\ldots,\delta_{k})$. We say that the curves defining $\Delta_{\mu}$ are parallel in $R$. Let $R^{\prime}=\tilde{\gamma}\setminus\mathrm{int}(R)$, where $\mathrm{int}(R)$ is the interior of $R$. There may be digon/triangle cells on $\mu$ in $R^{\prime}$ defined by the curves in $\Delta_{\mu}$. The next Lemma shows that we can get rid of such cells.
@@
We can apply a sequence of digon/triangle cell bypassing operations obtaining a modified sequence $\Delta^{\prime}_{\mu}$ of parallel curves in $R$ that satisfies the following conditions:
  1. (i)
    the resulting arrangement is as simple or simpler than $\Gamma$,
  2. (ii)
    there is no digon cell on $\mu$ in $R^{\prime}$ defined by a curve in $\Delta^{\prime}_{\mu}$ i.e., any digon cells on $\mu$ in $R^{\prime}$ is defined by a curve not in $\Delta^{\prime}_{\mu}$
  3. (iii)
    there is no triangle cell on $\mu$ in $R^{\prime}$ defined by two curves in $\Delta^{\prime}_{\mu}$,
  4. (iv)
    any triangle cell on $\mu$ in $R^{\prime}$ does not involve any of the curves in $\Delta^{\prime}_{\mu}$ except possibly the first or the last curve (but not both together).
  5. (v)
    no new digon/triangle cells are created on $\gamma$ in $\tilde{\gamma}$ apart from those that lie on $\mu$.
@@
If there were no digon/triangle cells to start with on $\mu$ in $R^{\prime}$, then there is nothing to do. If the curves in $\Delta_{\mu}$ do not intersect in $R^{\prime}$, then they do not form digon/triangle cells on $\mu$.
We repeat the following two steps while there is a digon cell on $\mu$ in $R^{\prime}$ defined by a curve in $\Delta_{\mu}$, or a triangle cell on $\mu$ in $R^{\prime}$ defined by two curves in $\Delta_{\mu}$. Each operation maintains the invariant that the curves are 2-intersecting, and the number of intersections among the curves in $\Delta_{\mu}$ that lie in $R^{\prime}$ decreases. Thus, the process stops after a finite number of steps. The operations also modify the sequence $\Delta_{\mu}$ to maintain the invariant that the curves in $\Delta_{\mu}$ are parallel in $R$.
If there is a digon cell $D$ on $\mu$ in $R^{\prime}$ bound by a curve $\delta_{i}\in\Delta_{\mu}$, then bypassing $D$ decreases the number of intersections in $R^{\prime}$. By Corollary 3, the arrangement remains 2-intersecting. The curve $\delta_{i}$ now forms a digon cell in $R$ with $\gamma$. We apply one more digon bypassing operation so that $\delta_{i}$ does not intersect $\gamma$. We now remove $\delta_{i}$ from $\Delta_{\mu}$. Thus, the modified $\Delta_{\mu}$ still corresponds to a sequence of parallel curves in $R$.
If $\delta_{i}$ and $\delta_{i+1}$ form a triangle cell $T$ on $\mu$ in $R^{\prime}$, let $p$ be the vertex of $T$ where $\delta_{i}$ and $\delta_{i+1}$ intersect. Then, bypassing this minimal triangle cell by $\mu$ decreases the number of intersections in (the modified) $R^{\prime}$, as $p$ does not lie in $R^{\prime}$ any more. We now obtain a triangle cell on $\gamma$ defined by $\delta_{i}$ and $\delta_{i+1}$. We again apply triangle bypassing so that the intersection point $p$ between $\delta_{i}$ and $\delta_{i+1}$ lies outside $\tilde{\gamma}$. The curves $\delta_{i}$ and $\delta_{i+1}$ remain parallel in $R$, but their order is swapped. Hence, we swap the order of $\delta_{i}$ and $\delta_{i+1}$ in $\Delta_{\mu}$ so that $\Delta_{\mu}$ is a sequence of parallel curves in $R$. By Corollary 3, the arrangement remains 2-intersecting. Hence, we obtain a 2-intersecting arrangement with fewer intersections in $R^{\prime}$.
In each digon/triangle cell bypassing operation on $\mu$ and $\gamma$, we modify $\Delta_{\mu}$ so that the curves remain parallel in $R$. Note that bypassing a digon or triangle cell on $\mu$ may increase the number of digon/triangle cells on $\mu$ in $R^{\prime}$. However, the number of intersections among the curves in $\Delta_{\mu}$ in $R^{\prime}$ keeps decreasing.
Thus, after a finite number of steps, we have a sequence $\Delta^{\prime}_{\mu}=(\delta^{\prime}_{1},\ldots,\delta^{\prime}_{\,k^{\prime}})$ of parallel curves in $R$ so that there is no digon cell on $\mu$ in $R^{\prime}$ formed by a curve in $\Delta^{\prime}_{\mu}$, and there is no triangle cell on $\mu$ in $R^{\prime}$ formed by two curves in $\Delta^{\prime}_{\mu}$.
The only digon cells on $\mu$ in $R^{\prime}$ can be formed by curves not in $\Delta^{\prime}_{\mu}$. Similarly, the only triangle cells formed on $\mu$ in $R^{\prime}$ are bounded by $\mu$, one of the curves in $\{\delta^{\prime}_{1},\delta^{\prime}_{k^{\prime}}\}$, and a curve not in $\Delta^{\prime}_{\mu}$. Note that there cannot be a triangle cell formed by $\mu$, $\delta^{\prime}_{j}$ (for some $j$ s.t. $12\leq j\leq k^{\prime}-$ 1) and a third curve not in $\Delta^{\prime}_{\mu}$ as that would violate the assumption that $\Delta^{\prime}_{\mu}$ is a set of parallel curves in $R$. Finally, note that the digon/triangle cell bypassing operations applied do not modify any cell not on $\mu$. Figures 16 show the modification of the curves in $\Delta_{\mu}$.
@@
Now, we have the tools to deal with the case when $C$ is a triangle cell. Suppose that $C_{\alpha}$ is a triangle cell $T$ that does not contain $P$, defined by the curves $\alpha,\gamma$ and $\delta_{1}$. Let $R$ be the digon defined by $\alpha$ and $\gamma$ that contains $T$. Without loss of generality let $a_{1},c_{1},d_{1}$ denote the three vertices of $T$, where $c_{1}$ and $d_{1}$ are the intersection points of $\delta_{1}$ with $\alpha$ and $\gamma$, respectively. Let $\Delta_{\alpha}=(\delta_{1},\ldots,\delta_{k})$ be a maximal sequence of parallel curves in $R$ starting with $\delta_{1}$. Let $d_{k}$ be the intersection point of $\delta_{k}$ and $\gamma$ in $R$. Since $\Gamma$ is a non-sweepable arrangement, either the arc of $\gamma$ between $d_{k}$ to $b_{2}$ does not bound a triangle cell, or the triangle cell contains $P$, and is therefore non-sweepable. Similarly, since $\Gamma$ is non-sweepable, the arc of $\gamma$ between $a_{1}$ and $b_{1}$ is intersected by a curve in $\Gamma\setminus\{\alpha,\beta,\gamma\}$, or is a triangle cell containing $P$. We first apply Lemma 7.6 with $\mu=\alpha$, $R$, and $\Delta_{\alpha}$. Let $\Delta_{\alpha}=(\delta^{\prime}_{1},\ldots,\delta^{\prime}_{\ell})$ denote the sequence of parallel curves after the application of Lemma 7.6. By Lemma 7.6, the possible sweepable triangle cells in the new arrangement are those in $\tilde{\gamma}$ that have $\delta^{\prime}_{1}$ or $\delta^{\prime}_{\ell}$ as a side, i.e., the cells containing reference points $x_{1},y_{1},x_{2}$ or $y_{2}$ (See Figure 17(a)). The cell containing reference point $y_{1}$ cannot be a sweepable cell it either contains $P$, or there is at least one more curve other than $\gamma$ defining this cell that does not intersect $\alpha$ in $\tilde{\gamma}$. Similarly, the cell containing the reference point $x_{2}$ cannot be a sweepable cell, as this would imply that the cell in $\Gamma$ containing $x_{2}$ was already sweepable. Therefore, only the cells containing reference points $x_{1}$ or $y_{2}$ can be sweepable triangle cells. Observe that $\alpha,\gamma$ and $\delta^{\prime}_{\ell}$ form a minimal triangle with base $\delta^{\prime}_{\ell}$. By Corollary 5.4, bypassing this minimal triangle yields a simpler non-sweepable arrangement of 2-intersecting curves such that the cell with reference point $x_{1}$ lies outside $\tilde{\gamma}$, and cell containing reference point $y_{2}$ becomes a 4-sided cell. See Figure 17(b). Thus, we obtain a simpler non-sweepable arrangement. If $C_{\beta}$ is a triangle cell, we treat it similarly. Thus, in all cases, we obtain a simpler non-sweepable arrangement, contradicting the assumption that $\Gamma$ was a simplest non-sweepable arrangement. Hence, we can apply one of the sweep operations, and by Corollary 3, each sweep operation preserves the 2-intersecting property of the curves.